OpenStudy (anonymous):

Simplify the expression. 6 + (14 – 6) • 3 + 2 • 3

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

ok, you need to use pemdas (i think thats what its called) basically the order of operations, you start with the parenthesis 14-6=8 then you do 8•3=24 then you have 6+24+2•3 you do 2•3=6, then just add everything together! 6+24+6=36! So your answer is 36!

OpenStudy (anonymous):

I am only 10

OpenStudy (anonymous):

I'm sorry, i don't know why it has all those boxes, here...

OpenStudy (anonymous):

Do you know order of operations?

OpenStudy (anonymous):

sorta

OpenStudy (anonymous):

Basically, when solving, you solve in this order, from left to right: Parenthesis, Exponents, Multiplication, Division, Addition, Subtraction.

OpenStudy (anonymous):

so the answer is 36

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

listen to the pemdas song that helped me before

OpenStudy (anonymous):

umm @AielandElsa ur only 10 years old?

OpenStudy (anonymous):

yes